# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v4

@@ III-B Asymptotic Performance Analysis
For the case of $\theta_{\mathrm{max}}>0$, the inequality $\mathcal{G}(R,\frac{1}{2},\theta_{\mathrm{max}})>2\bar{P}\xi\left[1-\operatorname{cos}\theta_{R}+\operatorname{sin}\theta_{\mathrm{max}}\left(\operatorname{ln}\frac{1+\operatorname{sin}\theta_{R}}{\operatorname{cos}\theta_{R}}+\operatorname{ln}\frac{1-\operatorname{sin}\theta_{\mathrm{max}}}{1-\operatorname{sin}^{2}\theta_{\mathrm{max}}}\right)\right]>\mathcal{G}(R,\frac{1}{2},0)$ always holds, indicating that the SNR’s lower bound of the RA system is always higher than that of the fixed-antenna system when $R_{\mathrm{lb}}>r\operatorname{tan}\theta_{\mathrm{max}}$. As the UPA size $N_{x}\Delta,N_{y}\Delta\to\infty$, the radii of the inscribed and circumscribed disks of the UPA region $N_{x}\Delta\times N_{y}\Delta$ go to infinity, i.e., $R_{\mathrm{lb}},R_{\mathrm{ub}}\to\infty$. Therefore, the lower/upper bounds given by Theorem 2 approach to the same limit due to the identical form of the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ as $R\to\infty$. Based on the above, we draw a conclusion that for the UPA setting, the RA system also achieves a higher asymptotic SNR than the fixed-antenna system, as in the previous ULA case.
@@ IV Multi-User Case under Multipath Channel
In this section, we consider the general multi-user and multipath channel setup, i.e., $K>1$ and $Q\geq 0$. Specifically, an AO algorithm and a two-stage algorithm are proposed to solve problem (P1) suboptimally, which offer different trade-offs between system performance and computational complexity.
@@ IV-A AO Algorithm
To overcome the challenges posed by the non-concavity of the objective function in (14a) and the intricate coupling between the receive beamforming vectors $\{\mathbf{v}_{k}\}$ and the RAs’ pointing vectors $\{\vec{\mathbf{f}}_{n}\}$, an AO algorithm is proposed to alternately optimize the receive beamforming and RAs’ pointing vectors in an iterative manner for the multi-user system.
1) Receive Beamforming Optimization: For a given RA’s pointing matrix $\mathbf{F}$, the channel from user $k$ to the BS modeled in (10) becomes fixed. Accordingly, problem (P1) reduces to (by simplifying $\mathbf{h}_{k}(\mathbf{F})$ to $\mathbf{h}_{k}$)
@@
[…]
The SINR in (36a) is a generalized Rayleigh quotient with respect to $\mathbf{v}_{k}$, and thus the receive SINR for each user can be maximized by the MMSE beamforming [43, 44]. Accordingly, the optimal solution to problem (P4) can be obtained as
@@
where $\mathbf{C}_{k}\triangleq\sum_{j\neq k}^{K}{\bar{P}_{j}\mathbf{h}_{j}\mathbf{h}_{j}^{H}}+\mathbf{I}_{N}$ is the interference-plus-noise covariance matrix. To reduce the dimension of matrix inversion from $N\times N$ to $(K-1)\times(K-1)$, by applying the Woodbury matrix identity, $\mathbf{C}_{k}^{-1}$ is equivalently expressed as
@@
2) RA Pointing Vector Optimization: For a given receive beamforming matrix $\mathbf{V}$, by introducing a slack optimization variable $\eta$ to denote the minimum SINR, problem (P1) can be written as
[…]
@@
$\alpha_{k,n}$ $\triangleq\frac{1}{r_{k,n}}\sqrt{\frac{AG_{0}}{4\pi}}e^{-j\frac{2\pi}{\lambda}r_{k,n}},$ (41a)
$\beta_{k,n,q}$ $\triangleq\frac{1}{d_{q,n}t_{k,q}}\frac{\sqrt{AG_{0}\sigma_{q}}}{4\pi}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+t_{k,q})+j\chi_{q}}.$ (41b)
@@
$\operatorname{ln}\left(\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}\right)\geq\operatorname{ln}(\eta)+\operatorname{ln}\left(\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1\right),$ (43)
which is still difficult to handle since $h_{k}(\vec{\mathbf{f}}_{n})$ in (40) is neither convex nor concave due to the complex coefficients $\{\alpha_{k,n}\}$ and $\{\beta_{k,n,q}\}$. To tackle this challenge, we adopt the SCA technique to approximate constraint (43) as a convex constraint and obtain a local optimal solution to problem (P6) in an iterative manner. Without loss of generality, we present the procedure of the $(i+1)$-th iteration and denote the solutions of $\mathbf{F}$ and $\eta$ obtained in the $i$-th iteration by $\mathbf{F}^{(i)}$ and $\eta^{(i)}$, respectively. By using the first-order Taylor expansion at $\{\vec{\mathbf{f}}_{n}^{(i)}\}$, $|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}$ and $\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1\right)$ in (43) can be respectively linearized as $\Lambda_{k}^{(i+1)}(\mathbf{F})$ and $\Gamma_{k}^{(i+1)}(\mathbf{F})$, shown at the top of the next page, where $\mathbf{h}_{k,n}^{{}^{\prime}}\triangleq\frac{\partial h_{k}(\vec{\mathbf{f}}_{n}^{(i)})}{\partial\vec{\mathbf{f}}_{n}^{(i)}}=\tilde{\alpha}_{k,n}\vec{\mathbf{u}}_{k,n}+\sum_{q=1}^{Q}{\tilde{\beta}_{k,n,q}\vec{\mathbf{c}}_{q,n}}$ with $\tilde{\alpha}_{k,n}\triangleq\alpha_{k,n}p((\vec{\mathbf{f}}_{n}^{(i)})^{T}\vec{\mathbf{u}}_{k,n})^{p-1}$ and $\tilde{\beta}_{k,n,q}\triangleq\beta_{k,n,q}p((\vec{\mathbf{f}}_{n}^{(i)})^{T}\vec{\mathbf{c}}_{q,n})^{p-1}$. Similarly, an upper bound for $\operatorname{ln}(\eta)$ is obtained as $\Xi^{(i+1)}(\eta)\triangleq\operatorname{ln}(\eta^{(i)})+\frac{\eta}{\eta^{(i)}}-1$ by using its first-order Taylor expansions at $\eta^{(i)}$. In this way, constraint (43) can be approximated by
@@
$\Gamma_{k}^{(i+1)}(\mathbf{F})\triangleq\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})|^{2}}+1\right)+\frac{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}\mathrm{Re}\left\{\left(\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})\right)^{\ast}\sum_{n=1}^{N}{v_{k,n}^{\ast}\left(\mathbf{h}_{j,n}^{{}^{\prime}}\right)^{T}(\vec{\mathbf{f}}_{n}-\vec{\mathbf{f}}_{n}^{(i)})}\right\}}}{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})|^{2}}+1}.$ (45)
@@
[…]
However, problem (P7) is still non-convex due to the unit-modulus constraint for $\vec{\mathbf{f}}_{n}$ in (14c). For convenience, we first relax the equality constraint (14c) as $\|\vec{\mathbf{f}}_{n}\|\leq 1$, yielding the following problem.
[…]
@@
1:  Input: Pointing vector $\mathbf{F}^{(0)}$, minimum receive SINR $\eta^{(0)}$, and threshold $\varepsilon>0$.
@@
7:  until $|\frac{\eta^{(i+1)}-\eta^{(i)}}{\eta^{(i)}}|\leq\varepsilon$.
@@
3) Overall Algorithm: Based on the results presented in the previous two subproblems, we propose the AO algorithm for problem (P1) by applying the block coordinate descent (BCD) method in Algorithm 1. Specifically, all optimization variables in the original problem (P1) are partitioned into two blocks, i.e., $\{\mathbf{V},\mathbf{F}\}$. Then, the receive beamforming $\mathbf{V}$ and RA’s pointing matrix $\mathbf{F}$ are alternately optimized, by calculating (37) and solving problem (P8), respectively, while keeping the other block of variables fixed. Furthermore, the obtained solution in each iteration is used as the input for the next iteration. The computational complexity of the proposed AO algorithm mainly depends on the matrix inversion for calculating the receive beamforming and the use of CVX solver to optimize the RA pointing matrix, which are performed in an iterative manner. Therefore, the complexity order of Algorithm 1 is $\mathcal{O}\left(L\left(KN^{3}+N^{3.5}\operatorname{ln}(1/\varepsilon)\right)\right)$, where $L$ and $\varepsilon$ denote the required iteration number and accuracy threshold for algorithm convergence, respectively.
@@
Note that Algorithm 1 solves the relaxed problem where the equality constraint $\|\vec{\mathbf{f}}_{n}\|=1$ in the original problem (P1) is relaxed to the inequality constraint $\|\vec{\mathbf{f}}_{n}\|\leq 1$. Thus, in the solution obtained by Algorithm 1, if the pointing vector $\vec{\mathbf{f}}_{n}$ is unit-modulus, i.e., the equality in (14c) holds, then the relaxation is tight and the obtained solution is feasible to problem (P1). Otherwise, the pointing vector needs to be reconstructed as a unit vector based on the solution obtained by Algorithm 1, i.e., $\vec{\mathbf{f}}_{n}^{\star}=\frac{\vec{\mathbf{f}}_{n}}{\|\vec{\mathbf{f}}_{n}\|}$.
@@ IV-B Two-Stage Algorithm
In this subsection, we propose another low-complexity algorithm, namely the two-stage algorithm, to solve problem (P1) without the need for iteration. Specifically, the pointing vectors of all RAs are optimized based on the semidefinite relaxation (SDR) technique in the first stage, and the corresponding beamforming vector is obtained by the ZF beamformer in the second stage.
As observed in Section IV-A, the difficulty in optimizing the pointing vectors mainly comes from the power function structure as shown in (40) and the fractional structure of the SINR. According to the law of power conservation, as the directivity factor $p$ increases, the maximum antenna gain $G_{0}$ in the boresight direction becomes larger and the antenna main lobe becomes narrower. Nevertheless, the variation in parameter $p$ does not change the relative magnitude relationship of radiation power in different directions. As such, we consider a typical value of $p$ to eliminate the power function structure. Specifically, for the cosine-square gain pattern with $p=1$, the multipath channel between user $k$ and RA $n$ modeled in (40) is expressed as the following linear combination form,
@@
where $\mathbf{m}_{k,n}\triangleq\alpha_{k,n}\vec{\mathbf{u}}_{k,n}+\sum_{q=1}^{Q}{\beta_{k,n,q}\vec{\mathbf{c}}_{q,n}}$.
@@
with $0\leq\rho_{\mathrm{ZF},k}\leq 1$ denoting the SNR loss factor caused by the cancellation of inter-user interference with ZF beamforming. According to (51), the resultant SNR for user $k$ based on ZF beamforming mainly depends on channel power gain $\|\bar{\mathbf{h}}_{k}\|^{2}$ when the SNR loss factor $\rho_{\mathrm{ZF},k}$ is given by a reasonable value.
1) First stage: Let $\mathbf{m}_{k}\triangleq[\mathbf{m}_{k,1}^{T},\mathbf{m}_{k,2}^{T},\dots,\mathbf{m}_{k,N}^{T}]^{T}\in\mathbb{C}^{3N\times 1}$ and $\mathbf{f}\triangleq[\vec{\mathbf{f}}_{1}^{T},\vec{\mathbf{f}}_{2}^{T},\dots,\vec{\mathbf{f}}_{N}^{T}]^{T}\in\mathbb{R}^{3N\times 1}$. Then, we formulate the following problem to optimize the pointing vectors.
[…]
@@
[…]
@@
Since problem (P10) may not lead to a rank-one solution of $\bar{\mathbf{F}}$, the optimal objective value of problem (P10) serves as an upper bound of problem (P9). Thus, a rank-one approximation on $\bar{\mathbf{F}}^{\star}$ should be executed as an additional step to construct a feasible solution to problem (P9). If $\bar{\mathbf{F}}^{\star}$ is rank-one, we have $\bar{\mathbf{F}}^{\star}=\mathbf{f}^{\star}\mathbf{f}^{\star T}$, and $\mathbf{f}^{\star}$ will be a feasible and optimal solution to problem (P9). On the other hand, if the rank of $\bar{\mathbf{F}}^{\star}$ is larger than one, we define $\bar{\mathbf{f}}=\sqrt{\lambda_{\mathrm{max}}}\bm{\upsilon}_{\mathrm{max}}$ with $\lambda_{\mathrm{max}}$ and $\bm{\upsilon}_{\mathrm{max}}$ denoting the maximum eigenvalue and its corresponding eigenvector obtained through eigenvalue decomposition of $\bar{\mathbf{F}}^{\star}$, respectively, as our candidate solution to problem (P9) since the best rank-one approximation to $\bar{\mathbf{F}}^{\star}$ is given by $\mathbf{f}^{\star}=\lambda_{\mathrm{max}}\bm{\upsilon}_{\mathrm{max}}\bm{\upsilon}_{\mathrm{max}}^{T}$ [46].
@@
The computational complexity of the proposed two-stage algorithm depends on the SDP and singular value decomposition in the first stage, and the calculation of the ZF beamforming in the second stage. Thus, the complexity order of the two-stage algorithm is given by $\mathcal{O}\left(KN^{3}+(3N)^{3}+(3N)^{3.5}\right)$. Since the two-stage algorithm only needs to solve problem (P10) and calculate the beamforming in (50) for one time, it requires lower computational complexity than the AO algorithm, especially when the number of antennas is large or the convergence accuracy of the AO algorithm is strict.
Remark 3: Although the two-stage algorithm may experience some performance loss by setting $p=1$, it has lower complexity without the need for iteration. In contrast, the AO algorithm, while having higher computational complexity, is applicable to any value of $p$. It is worth noting that both the AO and two-stage algorithms can also be applied to the single-user setup with arbitrary $Q$ scatterer clusters by replacing the MMSE/ZF receive beamforming of (37)/(50) with MRC beamforming. Nevertheless, for the single-user setup with $Q=0$, since the AO and two-stage algorithms can only achieve suboptimal solutions and result in higher computational complexity, they are much less efficient than the optimal closed-form solution derived in (20) of Section III-A. On the other hand, although the original problem can be transformed into a pointing vector optimization problem by substituting the MMSE receive beamforming of (37) into (13), it is found that the objective function becomes even more complicated due to the presence of matrix inversion, which is generally difficult to handle and thus not considered in this paper.

### v5

@@ III-B Asymptotic Performance Analysis
For the case of $\theta_{\mathrm{max}}>0$, the inequality $\mathcal{G}(R,\frac{1}{2},\theta_{\mathrm{max}})>2\bar{P}\xi\left[1-\operatorname{cos}\theta_{R}+\operatorname{sin}\theta_{\mathrm{max}}\left(\operatorname{ln}\frac{1+\operatorname{sin}\theta_{R}}{\operatorname{cos}\theta_{R}}+\operatorname{ln}\frac{1-\operatorname{sin}\theta_{\mathrm{max}}}{1-\operatorname{sin}^{2}\theta_{\mathrm{max}}}\right)\right]>\mathcal{G}(R,\frac{1}{2},0)$ always holds, indicating that the SNR’s lower bound of the RA system is always higher than that of the fixed-antenna system when $R_{\mathrm{lb}}>r\operatorname{tan}\theta_{\mathrm{max}}$. As the UPA size $N_{x}\Delta,N_{y}\Delta\to\infty$, the radii of the inscribed and circumscribed disks of the UPA region $N_{x}\Delta\times N_{y}\Delta$ go to infinity, i.e., $R_{\mathrm{lb}},R_{\mathrm{ub}}\to\infty$. Therefore, the lower/upper bounds given by Theorem 2 approach the same limit due to the identical form of the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ as $R\to\infty$. Based on the above, we conclude that for the UPA setting, the RA system also achieves a higher asymptotic SNR than the fixed-antenna system, as in the ULA case.
@@ IV Multi-User Case under Multipath Channel
In this section, we consider the general multi-user and multipath channel setup, i.e., $K>1$ and $Q\geq 0$. Specifically, an AO algorithm and a two-stage algorithm are proposed to solve problem (P1) suboptimally, which offer different trade-offs between system performance and complexity.
@@ IV-A AO Algorithm
To overcome the challenges posed by the non-concavity of the objective function in (14a) and the intricate coupling between the receive beamforming vectors $\{\mathbf{v}_{k}\}$ and the RA pointing vectors $\{\vec{\mathbf{f}}_{n}\}$, an AO algorithm is proposed to alternately optimize the receive beamforming and RA pointing vectors in an iterative manner for the multi-user system.
1) Receive Beamforming Optimization: For a given RA pointing matrix $\mathbf{F}$, the channel from user $k$ to the BS modeled in (10) becomes fixed. Accordingly, problem (P1) reduces to (by simplifying $\mathbf{h}_{k}(\mathbf{F})$ to $\mathbf{h}_{k}$)
$\text{(P4):}\quad\max_{\mathbf{V}}\quad$ $\min_{k}\quad\frac{\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}|^{2}}{\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}|^{2}}+1}$ (36a)
@@
The SINR in (36a) is a generalized Rayleigh quotient with respect to $\mathbf{v}_{k}$, and thus the receive SINR for each user can be maximized by the minimum mean-square error (MMSE) beamforming [47]. Accordingly, the optimal solution to problem (P4) can be obtained as
@@
where $\mathbf{C}_{k}\triangleq\sum_{j\neq k}^{K}{\bar{P}_{j}\mathbf{h}_{j}\mathbf{h}_{j}^{H}}+\mathbf{I}_{N}$ is the expected interference-plus-noise covariance matrix, which also corresponds to the sample covariance matrix over quasi-stationary channels. To reduce the dimension of matrix inversion from $N\times N$ to $(K-1)\times(K-1)$, by applying the Woodbury matrix identity, $\mathbf{C}_{k}^{-1}$ is equivalently expressed as
@@
2) RA Pointing Vector Optimization: For a given receive beamforming $\mathbf{V}$, by introducing a slack variable $\eta$ to denote the minimum SINR, problem (P1) can be written as
$\text{(P5):}\quad\max_{\eta,\mathbf{F}}\quad$ $\eta$ (39a)
@@
$\hskip-5.69046pt\alpha_{k,n}$ $\triangleq\frac{1}{r_{k,n}^{\mathrm{B-U}}}\sqrt{\frac{AG_{0}}{4\pi}}e^{-j\frac{2\pi}{\lambda}r_{k,n}^{\mathrm{B-U}}},$ (41a)
$\hskip-5.69046pt\beta_{k,n,q}$ $\triangleq\frac{1}{r_{q,n}^{\mathrm{B-C}}r_{k,q}^{\mathrm{U-C}}}\frac{\sqrt{AG_{0}\varsigma_{q}}}{4\pi}e^{-j\frac{2\pi}{\lambda}(r_{q,n}^{\mathrm{B-C}}+r_{k,q}^{\mathrm{U-C}})+j\chi_{q}}.$ (41b)
@@
$\hskip-2.84544pt\operatorname{ln}\hskip-2.84544pt\left(\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}\right)\hskip-2.84544pt\geq\hskip-2.84544pt\operatorname{ln}(\eta)\hskip-1.42271pt+\hskip-1.42271pt\operatorname{ln}\hskip-2.84544pt\left(\sum_{j\neq k}\hskip-2.84544pt{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}\hskip-2.84544pt+\hskip-2.84544pt1\right),\hskip-2.84544pt$ (43)
which is still difficult to handle since $h_{k}(\vec{\mathbf{f}}_{n})$ in (40) is neither convex nor concave due to the complex coefficients $\{\alpha_{k,n}\}$ and $\{\beta_{k,n,q}\}$. To tackle this challenge, we adopt the successive convex approximation (SCA) technique to approximate constraint (43) as a convex constraint and obtain a local optimal solution to problem (P6) in an iterative manner. Without loss of generality, we present the procedure of the $(i+1)$-th iteration and denote the solutions of $\mathbf{F}$ and $\eta$ obtained in the $i$-th iteration by $\mathbf{F}^{(i)}$ and $\eta^{(i)}$, respectively. By using the first-order Taylor expansion at $\{\vec{\mathbf{f}}_{n}^{(i)}\}$, $|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}$ and $\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1\right)$ in (43) can be respectively linearized as $\Lambda_{k}^{(i+1)}(\mathbf{F})$ and $\Gamma_{k}^{(i+1)}(\mathbf{F})$, shown at the top of the next page, where $\mathbf{h}_{k,n}^{{}^{\prime}}\triangleq\frac{\partial h_{k}(\vec{\mathbf{f}}_{n}^{(i)})}{\partial\vec{\mathbf{f}}_{n}^{(i)}}=\tilde{\alpha}_{k,n}\vec{\mathbf{u}}_{k,n}+\sum_{q=1}^{Q}{\tilde{\beta}_{k,n,q}\vec{\mathbf{c}}_{q,n}}$ with $\tilde{\alpha}_{k,n}\triangleq\alpha_{k,n}p\left[(\vec{\mathbf{f}}_{n}^{(i)})^{T}\vec{\mathbf{u}}_{k,n}\right]_{+}^{p-1}$ and $\tilde{\beta}_{k,n,q}\triangleq\beta_{k,n,q}p\left[(\vec{\mathbf{f}}_{n}^{(i)})^{T}\vec{\mathbf{c}}_{q,n}\right]_{+}^{p-1}$. Similarly, an upper bound for $\operatorname{ln}(\eta)$ is obtained as $\Xi^{(i+1)}(\eta)\triangleq\operatorname{ln}(\eta^{(i)})+\frac{\eta}{\eta^{(i)}}-1$ by using its first-order Taylor expansions at $\eta^{(i)}$. In this way, constraint (43) can be approximated by
@@
$\hskip-5.69046pt\Gamma_{k}^{(i+1)}(\mathbf{F})\triangleq\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})|^{2}}+1\right)+\frac{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}\mathrm{Re}\left\{\left(\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})\right)^{\ast}\sum_{n=1}^{N}{v_{k,n}^{\ast}\left(\mathbf{h}_{j,n}^{{}^{\prime}}\right)^{T}(\vec{\mathbf{f}}_{n}-\vec{\mathbf{f}}_{n}^{(i)})}\right\}}}{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})|^{2}}+1}.$ (45)
@@
However, problem (P7) is still non-convex due to the unit-norm constraint for $\vec{\mathbf{f}}_{n}$ in (14c). For convenience, we first relax the equality constraint (14c) as $\|\vec{\mathbf{f}}_{n}\|\leq 1$, yielding the following problem.
$\text{(P8):}\quad\max_{\eta,\mathbf{F}}\quad$ $\eta$ (48a)
@@
1: Input: Pointing vector $\mathbf{F}^{(0)}$, minimum receive SINR $\eta^{(0)}$, threshold $\varepsilon_{2}>0$, and maximum iteration number $I$.
@@
7:until $|\frac{\eta^{(i+1)}-\eta^{(i)}}{\eta^{(i)}}|\leq\varepsilon_{2}$ or $i>I$.
@@
3) Overall Algorithm: Based on the results presented in the previous two subproblems, we propose the AO algorithm for problem (P1) by applying the block coordinate descent (BCD) method in Algorithm 1. Specifically, all optimization variables in the original problem (P1) are partitioned into two blocks, i.e., $\{\mathbf{V},\mathbf{F}\}$. Then, the receive beamforming $\mathbf{V}$ and RA pointing matrix $\mathbf{F}$ are alternately optimized, by calculating (37) and solving problem (P8), respectively, while keeping the other block of variables fixed. Furthermore, the obtained solution in each iteration is used as the input for the next iteration. The computational complexity of the proposed AO algorithm is mainly due to the matrix inversion for calculating the receive beamforming and the use of CVX solver to optimize the RA pointing matrix, which are performed in an iterative manner. Therefore, the complexity order of Algorithm 1 is $\mathcal{O}\left(I\left(KN^{3}+N^{3.5}\operatorname{ln}(1/\varepsilon_{1})\right)\right)$, where $\varepsilon_{1}$ is the given solution accuracy and $I$ denotes the required iteration number for algorithm convergence.
@@
From a feasibility perspective, Algorithm 1 solves the relaxed problem where the equality constraint $\|\vec{\mathbf{f}}_{n}\|=1$ in the original problem (P1) is relaxed to the inequality constraint $\|\vec{\mathbf{f}}_{n}\|\leq 1$. Thus, in the solution obtained by Algorithm 1, if the pointing vector $\vec{\mathbf{f}}_{n}$ is unit-modulus, i.e., the equality in (14c) holds, then the relaxation is tight and the obtained solution is feasible to problem (P1). Otherwise, the pointing vector needs to be reconstructed as a unit vector based on the solution obtained by Algorithm 1, i.e., $\vec{\mathbf{f}}_{n}^{\star}=\frac{\vec{\mathbf{f}}_{n}}{\|\vec{\mathbf{f}}_{n}\|}$.
@@ IV-B Two-Stage Algorithm
In this subsection, we propose another low-complexity algorithm, namely the two-stage algorithm, to solve problem (P1) without the need for iteration. Specifically, the pointing vectors of all RAs are optimized using the semidefinite relaxation (SDR) technique in the first stage, and the corresponding beamforming vector is obtained by the ZF beamformer in the second stage.
As observed in Section IV-A, the main difficulty in optimizing the pointing vectors lies in the intricate power function structure as shown in (40) and the fractional structure of the SINR. According to the law of power conservation, as the directivity factor $p$ increases, the maximum antenna gain $G_{0}$ in the boresight direction becomes larger and the antenna main lobe becomes narrower. Nevertheless, the variation in parameter $p$ does not change the relative magnitude relationship of radiation power in different directions. As such, we consider a typical value of $p$ to eliminate the power function structure. Specifically, for the cosine-square gain pattern with $p=1$, the multipath channel between user $k$ and RA $n$ modeled in (40) is expressed as the following linear combination form,
@@
where $\mathbf{m}_{k,n}\triangleq\alpha_{k,n}\vec{\mathbf{u}}_{k,n}+\sum_{q=1}^{Q}{\beta_{k,n,q}\vec{\mathbf{c}}_{q,n}}$. For computational simplicity, the positive operators in (6) and (7) are omitted in (49) since all the users and scatterer clusters are confined to the front half-space of the BS.
@@
with $0\leq\rho_{\mathrm{ZF},k}\leq 1$ denoting the SNR loss caused by the cancellation of inter-user interference with ZF beamforming. According to (51), the resultant SNR for user $k$ based on ZF beamforming mainly depends on channel power gain $\|\bar{\mathbf{h}}_{k}\|^{2}$ when the SNR loss $\rho_{\mathrm{ZF},k}$ is given by a reasonable value.
1) First stage: Let $\mathbf{m}_{k}\triangleq[\mathbf{m}_{k,1}^{T},\mathbf{m}_{k,2}^{T},\dots,\mathbf{m}_{k,N}^{T}]^{T}\in\mathbb{C}^{3N\times 1}$ and $\mathbf{f}\triangleq[\vec{\mathbf{f}}_{1}^{T},\vec{\mathbf{f}}_{2}^{T},\dots,\vec{\mathbf{f}}_{N}^{T}]^{T}\in\mathbb{R}^{3N\times 1}$. We formulate the following problem to optimize the pointing vectors.
$\text{(P9):}\quad\max_{\omega,\mathbf{f}}\quad$ $\omega$ (53a)
@@
$\hskip-4.26773pt\text{(P10):}\quad\max_{\omega,\bar{\mathbf{F}}}\quad$ $\omega$ (54a)
@@
Since problem (P10) may not lead to a rank-one solution of $\bar{\mathbf{F}}$, the optimal objective value of problem (P10) serves as an upper bound of problem (P9). Thus, a rank-one approximation on $\bar{\mathbf{F}}^{\star}$ should be executed as an additional step to construct a feasible solution to problem (P9). If $\bar{\mathbf{F}}^{\star}$ is rank-one, we have $\bar{\mathbf{F}}^{\star}=\mathbf{f}^{\star}\mathbf{f}^{\star T}$, and $\mathbf{f}^{\star}$ will be a feasible and optimal solution to problem (P9). On the other hand, if the rank of $\bar{\mathbf{F}}^{\star}$ is larger than one, we define $\bar{\mathbf{f}}=\sqrt{\lambda_{\mathrm{max}}}\bm{\upsilon}_{\mathrm{max}}$, with $\lambda_{\mathrm{max}}$ and $\bm{\upsilon}_{\mathrm{max}}$ denoting the maximum eigenvalue and its corresponding eigenvector obtained through eigenvalue decomposition of $\bar{\mathbf{F}}^{\star}$, respectively. This serves as our candidate solution to problem (P9) since the best rank-one approximation to $\bar{\mathbf{F}}^{\star}$ is given by $\mathbf{f}^{\star}=\lambda_{\mathrm{max}}\bm{\upsilon}_{\mathrm{max}}\bm{\upsilon}_{\mathrm{max}}^{T}$ [18].
@@
The computational complexity of the proposed two-stage algorithm is mainly due to the SDP and singular value decomposition in the first stage, and the calculation of the ZF beamforming in the second stage. Thus, the complexity order of the two-stage algorithm is given by $\mathcal{O}\left(KN^{3}+(3N)^{3}+(3N)^{3.5}\right)$. Since the two-stage algorithm only needs to solve problem (P10) and calculate the beamforming in (50) for one time, it requires lower computational complexity than the AO algorithm.
Remark 2: Although the two-stage algorithm may experience some performance loss by setting $p=1$, it has lower complexity without the need for iteration. In contrast, the AO algorithm, while having higher computational complexity, is applicable to any value of $p$. It is worth noting that both the AO and two-stage algorithms can also be applied to the single-user setup with arbitrary $Q$ scatterer clusters by replacing the MMSE/ZF receive beamforming of (37)/(50) with MRC beamforming. Nevertheless, for the single-user setup with $Q=0$, since the AO and two-stage algorithms can only achieve suboptimal solutions and result in higher computational complexity, they are much less efficient than the optimal closed-form solution derived in (20) of Section III-A. On the other hand, although the original problem can be transformed into a pointing vector optimization problem by substituting the MMSE receive beamforming of (37) into (13), it is found that the objective function becomes even more complicated due to the presence of matrix inversion, which is difficult to handle and thus not considered in this paper.